OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. -x + 5y = 1 2x + 4y = -4 2(-5y + 1) + 4y = -4 2x + 4(-5y + 1) = -4 2x + 4(5y - 1) = -4 2(5y - 1) + 4y = -4

OpenStudy (anonymous):

-x+5y=1 5y=1+x 5y-1=x Therefore 2(5y-1)+4y=-4 is the equation

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

your welcome